OpenStudy (dannyo19):

The function f satisfies the equation f(x)+f(y)=f(x+y)-xy-1 for every pair x, y among all real numbers. If f(1)=1, find all integers n (if any) such that f(n)=n

OpenStudy (xapproachesinfinity):

i would think this way for a start f(x+y)=f(x)+f(y)+xy+1 f(1)=1 f(2)=4 So far i'm thinking of a sequence or something appearing just a feeling though

OpenStudy (dannyo19):

How do you suppose f(2)=4?

OpenStudy (xapproachesinfinity):

we are given f(1) i just made x=y in that case, but again I'm just trying to see anything useful

OpenStudy (xapproachesinfinity):

hmm let's see f(n)=n ==> f(n)+f(n)=f(2n)-n^2-1 2f(n)=f(2n)-n^2-1 f(2n)=n^2+2n+1=(n+1)^2

myininaya (myininaya):

maybe we can use f(1)=1 first to get f(x)+f(1-x)=-x(1-x) then use that f(n)=n where n is an integer to have x+(1-x)=-x(1-x) then solve this equation and see if we get x as an integer...

OpenStudy (xapproachesinfinity):

interesting @myininaya I haven't thought of using both, but i'm not in your league guys hehe

OpenStudy (xapproachesinfinity):

x^2-x-1=0 <== this has no integer solutions

imqwerty (imqwerty):

okay we have this-> \(f(1)=1\) now \(f(1)+f(0)=f(1+0)-1 \times 0 -1\) from here we get-> \(f(0)=-1\) now \(f(-1)+f(1)=f(-1+1)-1 \times 1 -1\) \(f(-1)+1=-1 -1 -1\) \(f(-1)=-4\) now for any general number say \(a\) \(f(a-1) + f(a+1)=f(2a) - (a-1)(a+1) -1\) ----> equation p now i do this- lets say \(n\) is the number for which \(f(n)=n\) so equation 1 equation 2 \(\color{blueviolet}{f(n)+f(1)=f(n+1)-n-1}\) \(\color{orange}{f(n)+f(-1)=f(n-1)+n-1}\) now add equation 2 and equation 1 we get this- \(2f(n) +f(1)+f(-1)=f(n-1)+f(n+1)-2\) now putting \(f(n)=n\) and using equation p we can write-> \(2n+1-4=f(2n) -(n-1)(n+1)-1\) now @xapproachesinfinity showed that \(f(2n)=(n+1)^2\) putting that in our equation we get- \(2n+1-4=(n+1)^2-n^2+1-1\) simplifying we get- \(2n-3=1+2n\) :) \(-3=1\) so no n possible o-0 except 1?

OpenStudy (xapproachesinfinity):

wow so the equation arrived by @myininaya is giving the same hint no solution except the given one f(1)=1

OpenStudy (qwertty123):

I think @imqwerty is getting it. Keep going.